James can produce either 25 widgets or 10 gadgets an hour, while John can produce either 10 widgets or 5 gadgets an hour. Which of the following statements is true?

Group of answer choices

A. John has an absolute advantage in producing both widgets and gadgets and thus has no reason to trade with James.

B. Since James has an absolute advantage in both goods, no mutual gains from trade are possible.

C. Mutually beneficial trade can occur if James trades widgets for gadgets with John

D. Mutually beneficial trade can occur if James trades gadgets for widgets with John
Answer #1

Here James has the absolute advantage in rpoduction of both the goods but it does not mean that there can not be a mutual beneficial trade. So to know the trade pattern calculate comparative advanatge.

James can produce either 25 widgets or 10 gadgets an hour:

Opportunity cost of producing 1 Widget= 10/25 or 0.4 Gadgets

Opportunity cost of producing 1 gadget= 25/10 or 2.5 widgets

John can produce either 10 widgets or 5 gadgets an hour:

Opportunity cost of producing 1 Widget= 5/10 or 0.5 Gadgets

Opportunity cost of producing 1 gadget= 10/5 or 2 widgets

James has the lower opportunity cost of producing widget so it will produce widget.

John has the lower opportunity cost of producing gadget so it will produce gadget.

Correct answer:

C. Mutually beneficial trade can occur if James trades widgets for gadgets with John